OpenStudy (anonymous):

Willie runs 4 miles in 48 minutes. If Willie runs at the same rate, how many miles can he run in 84 minutes? 7 5 11 9

OpenStudy (anonymous):

@iGreen do u know?

OpenStudy (anonymous):

I got it

OpenStudy (anonymous):

7

OpenStudy (igreen):

We can set up a proportion. \(\dfrac{4}{48} = \dfrac{x}{84}\)

OpenStudy (igreen):

Multiply 84 to both sides: \(\dfrac{4 \times 84}{48}\) Simplify that @ineedurhelpnow

OpenStudy (igreen):

@Kaity_bugg123 Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

you were correct @Kaity_bugg123

OpenStudy (anonymous):

84/12

OpenStudy (anonymous):

thanks you guys @iGreen and @Eliza002